OpenStudy (anonymous):

I am supposed to solve this by the elimination method 2x + 3y =1 and 4x + 6y =2 is there no soulution to this problem?

OpenStudy (anonymous):

multiply the first equation with 2 u will get the second equation

OpenStudy (anonymous):

2x + 3y = 1 --------(1) 4x + 6y = 2 --------(2) Multiply (1) by 2 and subtract (2) from the resulting equation 2(2x + 3y = 1) gives 4x + 6y = 2 which is same as (2) so their subtraction results in zero. These two equations represent the same line and hence have infinite solutions

OpenStudy (radar):

parallel lines that do not intersect for a solution .

OpenStudy (anonymous):

@radar both equations represent the same line and hence they have infinite solutions

OpenStudy (anonymous):

divide the second eqn by 2 and u get the first eqn !!!!

OpenStudy (radar):

y=-2/3 (x) +1/3 y=-2/3x + =1/3 You're right Harkirat, they are identities, same line.

OpenStudy (anonymous):

if a1/a2 = b1/b2 = c1/c2 then the lines coincide and have infinitely many solutions

OpenStudy (anonymous):

thx for the medal can u pls kindly become my fan, i need some more.......